OpenStudy (anonymous):

For questions 3-4, let X = {x|x is a whole number less than 15}, Y = {x|x is a multiple of 3 less than 15}, Z = {x|x is a real number greater than or equal to 5.5} 3. What is X *upside down U* Y? A) {0,3,6,9,12} B) {3,6,9,12,15} C) {…, -6,,3,0,3,6,9,…} D) {…, -6,-3,0,3,6,9,12} 4. What is X U Y ? A) {0,3,6,9,12} B) {0,1,2,3,4,5,6,7,8,9,10,11,12,13,14} C) {0,3,6,9,12} D) {-3,0,3,6,9,12}

OpenStudy (anonymous):

@MSxSammiBooxX

OpenStudy (anonymous):

for question 3 u mean

OpenStudy (anonymous):

yeah, and for 4, its that symbol but in the normal "U" shape

OpenStudy (anonymous):

well,u know \(X = (- \infty , 15 )\) and \(Y = \left\{ 0,3,6,9,12 \right\}\) now can u say what members are in both \(X\) and \(Y \)

OpenStudy (anonymous):

I can ?

OpenStudy (anonymous):

wait...in the question it is said "a whole number" it mean all real numbers?

OpenStudy (anonymous):

or integers?

OpenStudy (anonymous):

I don't know. What I posted is word for word what is says

OpenStudy (anonymous):

but there would be no prob for the question 3...can u say ?

OpenStudy (anonymous):

I don't know, thats why I'm asking

OpenStudy (anonymous):

i know,well i wrote X and Y for u , if we suppose Y contains integers or real numbers that would not make different (because X contains integer ) , u see the members that are both in X and Y are 0,3,6,9,12

OpenStudy (anonymous):

and . . . you're losing me :)

OpenStudy (anonymous):

didn't get?

OpenStudy (anonymous):

Ok. I get that you wrote X and Y for me BUT, I don't understand how that pertains to this particular question

OpenStudy (anonymous):

please explain ? :)

OpenStudy (anonymous):

u want the members which are common in both X and , right?

OpenStudy (anonymous):

X and Y, sure

OpenStudy (anonymous):

well,so the numbers that are both in X and Y are 0,3,6,9,12

OpenStudy (anonymous):

Got it

OpenStudy (anonymous):

So 3 is A ?

OpenStudy (anonymous):

yes ;)

OpenStudy (anonymous):

YAY !

OpenStudy (anonymous):

and for forth one i'm not sure of X but as the answers show,we should only consider integers so X = {...,-1,0,1,2,...,14 }

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

i don't know what does "whole number" mean, @mathmath333 what does that mean?

OpenStudy (anonymous):

i'm not sure

OpenStudy (mathmath333):

\(whole~number=\{0,1,2,2-----------\{\infty^{+}\}\)

OpenStudy (anonymous):

thanks,so B is correct ;)

OpenStudy (anonymous):

THANKS YOU GUYS !!!

OpenStudy (anonymous):

YAY ! They are all right ! @ _ @

OpenStudy (anonymous):

welcome ;)